# Constant-Size Certificates for Leader Election in Chordal Graphs and Related Classes

### v1

@@ Abstract
In distributed computing a certification scheme consists of a set of states and conditions over those states that enable each node of a graph to efficiently verify the correctness of a solution to a given problem. This work focuses on two fundamental problems: leader election and spanning tree construction. For each problem, we present a constant-size (per edge), local certification scheme, where the conditions available to each node can only refer to the graph induced by its one-hop neighborhood. In particular, we provide certification schemes for leader election in chordal and $K_{4}$-free dismantlable graphs and for spanning tree construction in dismantlable graphs, assuming a root is given. For chordal graphs, our leader election certification scheme additionally ensures an acyclic orientation, a property that is not generally verifiable using constant-size certificates in arbitrary graphs. To the best of our knowledge, these are the first local certification results tailored to these graph classes, potentially highlighting structural properties useful for verifying additional problems. Finally, we propose an algorithm that automatically transforms any certification scheme into a silent self-stabilizing algorithm (i.e., an algorithm that automatically recovers from faults) by adding only one extra state to the set of states of the certification scheme, assuming a Gouda fair scheduler. This transformation may be of independent interest.
@@
Local Certification, Distributed Verification, Proof Labelling Schemes, Chordal Graphs, Dismantlable Graphs, Leader Election, Distributed Computing
@@ 1 Introduction
Local certification [feuilloley2019introduction], also known as proof labelling schemes [korman2005proof], locally checkable proofs [goos2016locally] or distributed verification, verifies global properties through local conditions. Each node individually assesses these conditions and either accepts (i.e., all conditions are satisfied) or rejects (i.e., at least one condition is violated). An error is detected if any node rejects. For example, to verify that a graph is a cycle and not a path, each node accepts if its degree is two and rejects if its degree is one. However, the inverse problem of accepting a path and rejecting a cycle is not possible with certificates of constant size, even if the nodes are not identical [feuilloley2019introduction].
@@
The ability to continue functioning and recovering from faults is essential for distributed systems. Dijkstra [dijkstra1974stabilising] introduced the concept of self-stabilization as a generic way to model various faults and guarantee the correct behavior of an algorithm regardless of transient faults. An algorithm is self-stabilizing if it eventually converges to a correct output from any arbitrary initial configuration (see also [dolev2000book]). A self-stabilizing algorithm is called silent if when it reaches a correct configuration, it remains in that configuration unless a fault occurs. On the other hand, a self-stabilizing algorithm that is not silent can transition between correct configurations as the algorithm perpetually continues computing. The design of a self-stabilizing algorithm typically consists of three components: (a) Detect when the global configuration is faulty (b) Reach a non-faulty configuration (c) Solve the original task. The first aspect naturally connects self-stabilization with local certification, where nodes maintain local certificates that allow verification of the global configuration via local checks. For instance, in the self-stabilizing spanning tree algorithm of [chen1991spanning], each node stores a pointer to its parent (as a way to solve the task) and its distance from the root (as a way to verify the solution). Without the latter, cycles cannot be locally detected.
As a form of fault-tolerance, verification is not the main task to be solved within a system and as a result it should not consume too many resources. Hence, in this work, we only consider certificates of constant size per edge. A significant challenge in this setting is that in general problems such as the detection of cycles mentioned previously require more than constant space. However, the existence of models such as Population Protocols [aspnes2009population] and Stone Age Computing [emek2013stoneage] which only use constant space demonstrates that such constraints do not preclude the existence of solutions to non-trivial problems.
@@
The main problem that we consider in the following sections is Leader Election [LeLann77], which is an important primitive in distributed systems as it allows the coordination of processes in the system. The goal in this problem is for a unique node to enter and remain in the elected state indefinitely. A usual assumption within this context, that we do not make in this work, is that nodes have distinct identifiers. In the anonymous setting, the problem cannot be solved without additional assumptions due to symmetries [angluin1980local]. One way to overcome this impossibility is to characterize graph classes where the problem can be solved (e.g., as in [yamashita1996computing]) or to employ randomization (e.g., [itai1990symmetry]). In this work, we use the former method and provide certificates for leader election for the case of chordal and $K_{4}$-free dismantlable graphs.
@@
Both chordal graphs and dismantlable graphs can be characterized by an ordering of their vertices. Chordal graphs admit an ordering called a simplicial elimination scheme, in which each vertex is adjacent to a clique of nodes that appear later in the ordering. Similarly, dismantlable graphs admit a dismantling order, where each vertex $v$ has a neighbor that is adjacent to all other neighbors of $v$ appearing later in the ordering. Dismantlability, in particular, is a property shared by many important classes in metric graph theory, such as bridged [FaJa, SoCh] and Helly graphs [BP-absolute, BaPr]. In this work, we design certificates that take advantage of these orderings to define and locally verify an orientation of the edges incident to each node, ensuring that the graph contains a unique sink node, which we define to be the leader.
@@ 1.1 Related Work
In the LOCAL model introduced by Linial [linial1987local, linial1992journal], it is assumed that nodes exchange messages with neighboring nodes until a task is solved. The complexity of an algorithm within this model thus depends on the size of the messages and on the number of synchronous communication rounds. An alternative definition of the model that is more appropriate for our setting, is to assume that every node can see the state of each node in its neighborhood up to some distance $d$ and based on this partial knowledge of the graph, choose an output. The goal in this case is to minimize $d$ and the size of states for some given task. What can be computed within this setting has received significant attention (e.g., [vishkin1986determinisitic, goldberg1987parallel, luby1986simple, alon1986fast]). The concept of locally verifiable problems was formalized by Naor and Stockmeyer in their seminal paper [naor1993lcl, naor1995journal] in which they defined the family of Locally Checkable Labeling (LCL) problems. A problem is in LCL if given the view, consisting of constant-sized input and output labels, at some constant distance $d$ for each node, there exists an algorithm that can decide whether a computed solution is correct. Although LCL problems were continually explored since their introduction, no groundbreaking results emerged until 2016 when Brandt et al. [brandt2016lower] led to a series of papers (e.g., [balliu2018classes, balliu2018locally, balliu2020randomness]) that gradually provided a complete picture on the classification of the round complexity of all LCL problems in four broad classes [suomela2020landscape].
@@
Recently, Bousquet et al. [feuilloley2025certification] proposed a different direction of classification based on space instead of round complexity. Several results are given in [feuilloley2025certification], characterizing both “gaps” (i.e., pairs of functions $f(n),g(n)$, with $f(n)<g(n)$ where no property has optimal certificate size in $(f(n),g(n))$ for a given class of graphs) as well as proving the existence of properties with optimal certificate sizes for various classes of graphs including paths, cycles and caterpillars. In this work we consider chordal and ($K_{4}$-free) dismantlable graphs which are not explicitly studied in [feuilloley2025certification]. Our approach is complementary to that of [feuilloley2025certification] in the sense that we only consider specific problems but for larger classes of graphs.
The most similar paper to our work is [chalopin2024stabilising]. Therein, the authors propose a certificate and a set of conditions with constant memory per node, ensuring a unique sink for a simply connected subset of the regular triangular grid, where simply connected means that the configuration does not contain holes. Observe that simply connected parts of triangular grids are $K_{4}$-free dismantlable graphs, hence this work (in particular Section˜4) generalizes the result of [chalopin2024stabilising] and extends this idea to additional classes of graphs. Additionally, our definition of view is similar to the concept of binoculars introduced in [chalopin2015binoculars]. Binoculars provide each node $v$ with all information on the graph induced by the neighborhood of $v$ at distance one.
@@ 1.2 Our Contributions
We define constant-size certification schemes for verifying the existence of a unique sink in chordal graphs and $K_{4}$-free dismantlable graphs, under the assumption that each node can see the subgraph induced by its one-hop neighbors and the certificates of its neighbors. This notion of view can be seen as an intermediate assumption, requiring more information than standard radius one visibility but without accessing nodes at distance two. For chordal graphs, the certification scheme also ensures a valid acyclic orientation, which is particularly important as acyclicity cannot, in general, be verified with constant-size certificates. In the case of dismantlable graphs, we also provide a certification scheme to ensure the existence of a spanning tree, assuming there is a distinct node marked as the root. The certification scheme verifying the existence of a unique sink in $K_{4}$-free dismantlable graphs can be used to verify the existence of such a root. All of our certificates require only constant size per edge.
In the last section, we show that given a constant-size certificate, a silent self-stabilizing algorithm can always be constructed with constant memory per process, assuming a Gouda fair scheduler. Let $k$ be the number of bits needed to verify a task $T$ using a certificate. Our algorithm transforms the certificate to a self stabilizing algorithm using $k+1$ bits. This is less than prior general transformations, such as the one in [BlinFP14], which requires $O(k+\log n)$ bits, where $n$ is the number of computational entities in the system, but works even under an unfair scheduler. As a result, Algorithm 1 can be used in systems with constant memory where [BlinFP14] cannot be employed.
Our results demonstrate that self-stabilization with strong global guarantees (i.e., unique sink, acyclic orientation) is possible with limited local memory in natural graph classes. That is, we show that combining certificate-based verification and Gouda fairness provides a way for designing fault-tolerant systems even when only constant memory is available.
@@ 2.1 Model
Let $G=(V,E,\delta)$ be a graph with vertex set $V$, edge set $E$ and a port labeling function $\delta$. For any vertex $v\in V$, let $N(v)$ denote its neighborhood and let $N[v]=N(v)\cup\{v\}$ denote its closed neighborhood. The 1-ball at $v$, denoted by $B_{1}(v)$, is the graph induced by $N[v]$. When $G$ is oriented, we write $N^{+}(v)$ (resp. $N^{-}(v)$) for the outgoing (resp. incoming) neighborhood of $v$ in $G$ and refer to nodes in $N^{+}(v)$ as outneighbors of $v$. We define the port labeling function to be $\delta=\{\delta_{v}\}_{v\in V}$, where $\delta_{v}:N(v)\rightarrow\{1,2,\ldots,deg(v)\}$ assigns a unique port number to each edge incident to $v$ for every $v\in V$ (i.e., $\delta_{v}$ is a bijection). We write $K_{4}$ to denote a complete graph with four vertices. A graph is $K_{4}$-free if it contains no subgraph that is isomorphic to a $K_{4}$. We write directed triangle to refer to a cycle of size three.
@@
[…]
  • The set of directed edges $\{\vec{vw}\text{ }|\text{ }v\neq r\}$, where $w=\emph{parent}_{G}(v)$, is a directed spanning tree.
We slightly adapt the definition of local certification of [feuilloley2025certification] to our setting. We assume that the nodes of $G$ do not have identifiers, that is, $G$ is anonymous. Let $L$ be a possibly empty set of input labels. If $L\neq\emptyset$, we say that $G$ is labeled and we call the mapping from $V$ to $L$ the input function. Let $C$ be a non-empty set of states, called the certificate. A certificate assignment on $G$ is a mapping $x_{C}:V\rightarrow C$. For a given certificate assignment $x_{C}$ and a node $v\in V$, the view of $v$ consists of the 1-ball $B_{1}(v)$ and the restriction of the port numbering function $\delta$, the input function (if defined) and the certificate assignment $x_{C}$ to $B_{1}(v)$. An example of the view of a node is given in Figure˜1.
Figure 1: Right: A graph with port labels (red) where each node is assigned a certificate in $\{0,1\}^{*}$ (black). Left: The view of the node marked in gray.
Let $R$ be a set of configurations of the form $(B,x_{C})$ where $B$ is a ball of radius 1 and $x_{C}$ is a certificate assignment defined on the nodes of $B$. We refer to $R$ as a set of local configurations. A local configuration is said to be correct with respect to $(C,R)$ if for every $v\in V$ the pair $(B_{1}(v),x_{C}|_{N[v]})$ belongs to $R$, where $x_{C}|_{N[v]}$ denotes the restriction of $x_{C}$ to $N[v]$. In other words, the local configuration is correct if the labeling in each ball of radius 1 satisfies the local conditions defining a correct configuration $R$. We refer to the pair $(C,R)$ as a certification scheme.
A certification scheme is used to solve a given task $\mathcal{T}=(O,P)$. The key property we prove, is that if a configuration is locally correct with respect to the certification scheme, then it is globally correct (i.e., it is a correct configuration for the task). Formally, let $out:C\rightarrow O$. We say that there exists a certification scheme $(C,R)$ for some task $(O,P)$ in a family of possibly labelled graphs $\mathcal{F}$ if for every $G\in\mathcal{F}$ the following two properties are satisfied [^1] The symbol $\circ$ denotes the composition operator.:
@@
$\forall G\in\mathcal{F},\text{ }\exists(G,x_{C})\text{ s.t. }\forall v\in V,\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R$
@@
$\forall(G,x_{c})\text{ s.t. }G\in\mathcal{F},\text{ if }\forall v\text{ }(B_{1}(v),x_{C}|_{N[v]})\in R\text{ then }(G,\emph{out}\circ x_{C})\in P$
In most of our certification schemes, the certificates correspond to an edge orientation. For these certificates, we denote the configuration as $\vec{G}$ and we adopt the following encoding for the certificates: the certificate assigned to each node $v\in V$ is composed of $\deg(v)$ bits where each bit encodes the direction of the corresponding incident edge. We assume that 0 denotes an incoming edge to $v$ and 1 denotes an outgoing edge from $v$. The bit at position $i\in\{0,\ldots,\deg(v)-1\}$ of the certificate corresponds to the direction of the edge at port $i$ of $v$. A visualization of this for the graph in Figure˜1 is given in Figure˜2.
Figure 2: The $i$-th bit of the certificate (black numbers) of each node $v$ corresponds to the direction of the edge incident to port $i$ (red numbers) of $v$, for $i\in\{0,\ldots,deg(v)-1\}$ where 0 denotes an incoming and 1 an outgoing edge. On the left we show the direction of the edges and on the right we additionally show the bit of the certificate of each node corresponding to each edge.

### v2

# Leader Election via Unique Sink Orientation

@@ Abstract
A Locally Checkable Labeling (LCL) is a specification describing a set of labels that are valid with respect to a set of conditions that characterize a local part of a solution to a global problem. Conditions can only refer to nodes and labels within a constant radius neighborhood of each node. This work studies local labeling schemes whose global consistency implies solutions to two classical problems: leader election and spanning tree construction. For each problem, we present a local labeling scheme using one bit per edge or equivalently $O(\Delta)$ bits per node (where $\Delta$ is the maximum degree in the graph), with conditions checkable within the graph induced by the one neighborhood of each node. For leader election, we show that global satisfaction of the conditions implies the existence of a unique sink in the graph, which we define to be a leader, while in the spanning tree setting it implies that a specific subset of edges induces a spanning tree rooted at a given node. We show those implications for $K_{4}$-free dismantlable and chordal graphs in the former case and for dismantlable graphs in the latter, assuming a root is given. For chordal graphs, the labeling implying a unique sink additionally induces an acyclic orientation. This property is not generally locally verifiable with constant-size labels in arbitrary graphs. To the best of our knowledge, these are the first local labeling results tailored to ($K_{4}$-free) dismantlable graphs, potentially highlighting structural properties useful for designing LCLs for additional problems. Finally, we present a generic transformation that converts any local labeling scheme into a silent self-stabilizing algorithm by adding only one extra state, assuming a Gouda fair scheduler. This transformation may be of independent interest.
@@
Locally Checkable Labeling, LCL, Leader Election, Dismantlable Graphs, Chordal Graphs, Self-Stabilization, Distributed Computing
@@ 1 Introduction
A fundamental notion in distributed computing is locality, which studies the extent to which global problems admit solutions based solely on local information. In a local algorithm, each node in a network must determine its part of a global solution based solely on the topology and states of its radius $r$ neighborhood, for some $r$ that is smaller than the size of the network. The challenge of such local algorithms is ensuring that these independently computed local solutions remain globally consistent and, as a result, solve a global task without the benefit of global visibility. Locality is concerned with two notions:
  1. 1.
    Solving a global problem: If a set of local conditions are satisfied then a global property holds.
  2. 2.
    Verifying the solution to a global problem: If a global property holds then a set of local conditions are satisfied.
@@
The ability to continue functioning and recovering from faults is essential for distributed systems. Dijkstra introduced the concept of self-stabilization [dijkstra1974stabilising] as a generic way to model various faults and guarantee the correct behavior of an algorithm regardless of transient faults. An algorithm is self-stabilizing if it eventually converges to a correct output from any arbitrary initial configuration (see also [dolev2000book]). A self-stabilizing algorithm is called silent [dolev1999silent] if when it reaches a correct configuration, it remains in that configuration unless a fault occurs. On the other hand, a self-stabilizing algorithm that is not silent can transition between correct configurations as the algorithm perpetually continues computing. The design of a self-stabilizing algorithm typically consists of three components: (a) Detect when the global configuration is faulty (b) Reach a non-faulty configuration (c) Solve the original task. Thus self-stabilization is naturally connected to the concept of locality as observed for the first time in [afek1997local]. In a local algorithm each node maintains: (i) a label which is the local part of a global solution and (ii) a certificate which is additional information that allows the verification of the global configuration via local checks. For instance, in the self-stabilizing spanning tree algorithm of [chen1991spanning], each node stores a label which is a pointer to its parent (as a way to solve the task) and a certificate which is its distance from the root (as a way to verify the solution). Without the latter, cycles cannot be locally detected. As shown initially by [naor1993lcl], there exist distributed problems that admit local solutions (i.e., solutions that can be verified without additional information) for which a certificate is not needed. The specification of the set of allowed labelings (i.e., local solutions) to these problems up to some constant radius $r$ are called Locally Checkable Labelings (LCL). An example of a distributed problem with a local solution is vertex coloring in which every node knows whether the solution is locally correct by only comparing its label (i.e., its color) to that of its neighbors at distance one.
The labels we define in this work are an orientation of the edges incident to each node. For each class of graphs we consider, we define a set of local conditions and we show that if all conditions we define for each family of graphs hold at every node, the global configuration contains a unique sink that we define to be a leader. Then we use this description of what a solution looks like locally to construct a silent self-stabilizing algorithm for leader election in specific classes of graphs that works under a Gouda fair scheduler [Gouda2001theory] which is the strongest kind of fairness [dubois2011taxonomy]. Our self-stabilizing algorithm is generic and works for any problem for which the solution can be defined as a local labeling. Results in a similar flavor that combine local checks [afek1997local] with a global “reset” [awerbuch1994reset] to some predetermined state in order to achieve self stabilization already exist [awerbuch1994stabilization]. However, contrary to algorithms such as [awerbuch1994reset, awerbuch1994stabilization, BlinFP14], our algorithm only needs one additional bit of memory to transform a labeling scheme to a self-stabilizing algorithm, due to Gouda fairness.
@@
Leader Election [LeLann77], is an important primitive in distributed systems as it allows the coordination of processes in the system. The goal in this problem is for a unique node to enter and remain in the distinct elected state indefinitely. A usual assumption within this context, that we do not make in this work, is that each node has a unique identifier. In the anonymous setting, the problem cannot be solved without additional assumptions due to symmetries [angluin1980local]. One way to overcome this impossibility is to characterize graph classes where the problem can be solved (e.g., as in [yamashita1996computing]) or to employ randomization (e.g., [itai1990symmetry]). In this work, we adopt the former approach and study the problem in specific classes of graphs. For each class of graphs we define an LCL problem and we show that any labeling satisfying the conditions is a solution to the leader election problem. We give a formal definition of the problem for each class of graphs and each set of conditions in Sections 3.1, 3.2 and 4.
@@
We focus on $K_{4}$-free dismantlable graphs. Dismantlable graphs are exactly the class of cop-win graphs in the classic Cops and Robbers game [aigner1984cops, nowakowski1983vertex, quilliot1978jeux] and include well-structured families such as bridged [FaJa, SoCh] and Helly [BP-absolute, BaPr] graphs, which are known for their convexity and metric properties. We briefly consider two additional classes: trees (Section 3.1) and chordal graphs (Section 3.2). Trees are particular chordal graphs, which in turn are particular dismantlable graphs. Both dismantlable and chordal graphs can be characterized by an ordering of their vertices. Chordal graphs admit an ordering called a simplicial elimination scheme, in which each vertex is adjacent to a clique of nodes that appear later in the ordering. Similarly, dismantlable graphs admit a dismantling order, where each vertex $v$ has a neighbor that is adjacent to all other neighbors of $v$ appearing later in the ordering.
@@ 1.1 Related Work
The concept of locality in distributed computing was formalized by Linial [linial1987local, linial1992journal] in the LOCAL model as a way to study how local computations can synthesize a solution to a global problem. The focus of this model is on the distance from which each node must gather information in order to compute its part of the global solution. Hence the complexity of an algorithm in LOCAL depends on the number of synchronous communication rounds. What can be computed within this setting immediately received significant attention (e.g., [vishkin1986determinisitic, goldberg1987parallel, luby1986simple, alon1986fast]) and is still an active area of research. A closely related model, CONGEST [peleg2000distributed], is defined in the same way as LOCAL with the restriction that nodes can only exchange bounded information in each round, typically logarithmic to the size of the network.
@@
In the same spirit, Naor and Stockmeyer defined the class of Locally Checkable Labeling (LCL) problems in their seminal paper [naor1993lcl, naor1995journal]. An LCL problem is specified by a constant radius $r$ and a finite set of allowed labeled radius $r$ neighborhoods such that, in graphs of bounded degree, a labeling is globally correct if and only if every node’s radius $r$ neighborhood belongs to this set. Intuitively the definitions of LOCAL and LCL are connected. This connection is made explicit in Chang’s [chang2024lcl] definition of LCL: “A distributed problem on bounded-degree graphs is a locally checkable labeling (LCL) if there is some constant $r$ such that the correctness of a solution can be checked locally in $r$ rounds of communication in the LOCAL model”. A systematic study of LCL problems was started in Brandt et al. [brandt2016lower] which then led to a series of papers (e.g., [balliu2018classes, balliu2018locally, balliu2020randomness]) that eventually formed a complete picture on the classification of the round complexity of all LCL problems in four broad classes [suomela2020landscape]. Recently, Bousquet et al. [feuilloley2025certification] proposed a different direction of classification of LCL problems based on space instead of round complexity. Several results are given in [feuilloley2025certification], characterizing both “gaps” (i.e., pairs of functions $f(n),g(n)$, with $f(n)<g(n)$ where no property has optimal certificate size in $(f(n),g(n))$ for a given class of graphs) as well as proving the existence of properties with optimal certificate sizes for various classes of graphs including paths, cycles and caterpillars. In this work we consider chordal and ($K_{4}$-free) dismantlable graphs which are not explicitly studied in [feuilloley2025certification]. Our approach is complementary to that of [feuilloley2025certification] in the sense that we only consider specific problems but for larger classes of graphs.
The most similar paper to our work is [chalopin2024stabilising], which also locally checks a set of conditions that ensure the existence of a unique sink defined to be a leader in a specific class of graphs. The main similarity between our work and [chalopin2024stabilising] is the assumption that nodes are anonymous. In [chalopin2024stabilising], the authors propose a constant-memory labeling and a set of sufficient conditions to ensure a unique sink for a simply connected subset of the regular triangular grid $S$, where simply connected means that the configuration does not contain “holes” (i.e., subsets of the triangular grid that do not belong in $S$ and are disconnected to each other). As simply connected parts of triangular grids are $K_{4}$-free dismantlable graphs, this work (in particular Section˜4) generalizes the result of [chalopin2024stabilising] and extends this idea to additional classes of graphs. The natural generalization of [chalopin2024stabilising] would be considering graphs with a contractible clique-complex, but this class is difficult to handle algorithmically, as recognizing these graphs is undecidable (this follows from a result of Novikov, as explained in [Tancer2016collapsible]). Dismantlable graphs form a class of graphs that have contractible clique-complexes, but they are easier to handle as they can be recognized in polynomial time [nowakowski1983vertex, quilliot1978jeux], making them a natural class to consider. Chordal graphs form a subclass of dismantlable graphs that is already very large as they contain all split graphs and there are $2^{\Theta(n^{2})}$ split graphs [BRW-split1985, oeis-split]. Observe that adding a universal vertex (i.e., a vertex neighboring to every other vertex in the graph) to any $n$-node graph results in an $(n+1)$-node dismantlable graph implying that the number of dismantlable graphs with $n+1$ vertices is larger than the number of $n$-node graphs. If the initial graph is additionally triangle-free, then the resulting graph is $K_{4}$-free and thus the number of $K_{4}$-free dismantlable graphs of size $n$ is $2^{\Theta(n^{2})}$ [EKR-KnFreeGraphs1976, oeis-triangle-free]. In fact, the class of $K_{4}$-free bridged graphs (which is a subclass of $K_{4}$-free dismantlable graphs is already very general. Indeed, Weetman [Weetman1994] shows that starting from any regular graph $G$ where every cycle has length at least 6, one can construct an infinite graph $H$ where the neighborhood of any vertex induces a graph isomorphic to $G$. This graph $H$ is an infinite $K_{4}$-free bridged graph and any ball of $H$ (and more generally any finite convex subgraph of $H$) induces a finite $K_{4}$-bridged graph.
Recently, [feuilloley2025proving] studied the verification aspect of a unique leader existing in anonymous grids and chordal graphs, giving an $O(\log D)$ distance based certification scheme, where $D$ is the diameter of the network. After the first version of the current paper appeared, the idea of [feuilloley2025proving] to verify the existence of a unique leader by a certificate encoding the distance of each node to the leader was extended to more classes of graphs in [ChChKo-meshed]. Notably, [ChChKo-meshed] verifies the existence of a unique leader in meshed graphs using certificates of constant size, where the certificate of each node is the modulo 3 distance to the leader. Meshed graphs is a class of graphs that also contains chordal, bridged, and Helly graphs. However, the classes of meshed and ($K_{4}$-free) dismantlable graphs are not comparable. In fact, meshed graphs are defined by a distance condition whereas dismantlable graphs do not admit such a metric characterization. As a result it is not clear whether a similar distance based labeling scheme can be defined for ($K_{4}$-free) dismantlable graphs. In this work, the labels we assign at each node encode an orientation that implies the existence of a unique sink, defined to be a leader. Hence, the size of our labels depends on the maximum degree of the graph. This means that we only obtain a constant size labeling in graphs of bounded degree. This work as well as the work of [feuilloley2025proving] and [ChChKo-meshed] assume that nodes have access to information that fits within the radius two verification rules. To the best of our knowledge no other papers consider leader election in LCL or related models in anonymous graphs.
@@ 1.2 Our Contributions
As already defined, an LCL is a specification describing a set of allowed radius $r$ neighborhoods, where “allowed” refers to conditions proving that the labels solve a global problem. An algorithm is then needed as a procedure that produces those labels. In this work:
1. We define a set of labels, which are usually edge orientations, and a set of conditions that define the allowed labelings within the graph induced by the one neighborhood for each node in $K_{4}$-free dismantlable graphs and related classes. The problem we solve is ensuring that the orientation of all edges incident to every node satisfies a number of conditions that we separately define for each class of graphs. We show that when those conditions globally hold, the network contains a unique sink that we define to be a leader. The size of the label at each node is deg bits (where deg is the degree of the node). As most of our labeling schemes are an orientation of all edges incident to each node, our approach results in constant size labels for graphs of bounded degree or bounded degeneracy [feuilloley2021degeneracy, feuilloley2023genus]. Although in the distributed verification and LCL literature it is usually assumed that nodes have unique identifiers, here we only consider anonymous nodes. Instead of the usual approach of using structural properties for the construction of the labels, in this work we capture the geometry of graphs in the labels. This geometric approach can be useful in practical geometric settings, as for example our result holds for planar triangulations where inner nodes have degree at least six. As planar graphs are also 5-degenerate, this is an example of a class of graphs with bounded degeneracy where our approach leads to constant-size labels.
Despite dismantlable graphs (and chordal graphs in Section 3.2) admitting an ordering of their vertices, the edge orientation satisfying our conditions is much more general as we need to show that every labeling satisfying the conditions results in an orientation with a unique sink. In fact, our LCL ensures a unique sink despite the orientation of edges allowing large cycles. This highlights the fact that with our orientation one cannot use the well known trick of constructing a rooted spanning tree and electing the root. In the particular subcase of chordal graphs, our labeling scheme additionally ensures an acyclic orientation. For the case of dismantlable graphs, we also provide a labeling scheme to construct a spanning tree, assuming a distinct node marked as the root is given as input. The LCL determining a unique sink in $K_{4}$-free dismantlable graphs can be used to calculate such a root.
2. We show that given an LCL, a silent self-stabilizing algorithm can always be constructed with one additional bit of memory per node, assuming a Gouda fair scheduler. Although [goos2016locally] already mentions that locally checkable proofs, which include LCLs, can be interpreted as nondeterministic local algorithms, here we formalize this intuition and extend it to defining a self-stabilizing algorithm. The one additional bit needed by Algorithm 1 to transform any $k$-bit labeling scheme to a self-stabilizing algorithm is less than prior general transformations, such as the one in [BlinFP14], which requires $O(k+\log n)$ bits, where $n$ is the number of computational entities in the system, but works even under an unfair scheduler. So, Algorithm 1 can be used in systems with constant memory where [BlinFP14] cannot be employed. In our case, this means that Algorithm 1 produces an edge orientation with a unique sink, for any initially arbitrarily directed (or undirected) $K_{4}$-free dismantlable or chordal graph.
@@ 2.1 Model
Let $G=(V,E,\delta)$ be a graph with vertex set $V$, edge set $E$ and a port labeling function $\delta$. For any vertex $v\in V$, let $N(v)$ denote its neighborhood and let $N[v]=N(v)\cup\{v\}$ denote its closed neighborhood. The degree of a vertex $v$ is the size of $N(v)$ and is denoted by $\deg(v)$. The 1-ball at $v$, denoted by $B_{1}(v)$, is the graph induced by $N[v]$. When $G$ is oriented, we write $N^{+}(v)$ (resp. $N^{-}(v)$) for the outgoing (resp. incoming) neighborhood of $v$ in $G$ and refer to nodes in $N^{+}(v)$ as outneighbors of $v$. We define the port labeling function to be $\delta=\{\delta_{v}\}_{v\in V}$, where $\delta_{v}:N(v)\rightarrow\{1,2,\ldots,\deg(v)\}$ assigns a unique port number to each edge incident to $v$ for every $v\in V$ (i.e., $\delta_{v}$ is a bijection). We write $K_{4}$ to denote a complete graph with four vertices. A graph is $K_{4}$-free if it contains no subgraph that is isomorphic to a $K_{4}$. We write directed triangle to refer to a directed cycle of size three.
@@
Let $O=\mathbb{N}$. Let $\operatorname{parent}_{G}(v)$ be the unique vertex $w\in N(v)$ such that $\delta_{v}(w)=x_{O}(v)$ if $1\leq x_{O}(v)\leq\deg(v)$ and let $\operatorname{parent}_{G}(v)=\bot$ otherwise (e.g., if $x_{O}(v)=0$). The set of correct configurations $P$ is the set of configurations $(G,x_{O})$ such that:
  • $\forall v\in V$, $0\leq x_{O}(v)\leq\deg(v)$,
  • ! $\exists!\text{ }r\in V$ such that $\operatorname{parent}_{G}(r)=\bot$ (i.e., such that $x_{O}(v)=0$),
  • The set of directed edges $\{\vec{vw}\mid v\neq r\text{ and }w=\operatorname{parent}_{G}(v)\}$ is a directed spanning tree.
We assume that the nodes of $G$ do not have identifiers, that is, $G$ is anonymous. Let $I$ be a possibly empty set of inputs and call the mapping from $V$ to $I$ the input function. Let $L$ be a non-empty set of states, called the labels. A labeling on $G$ is a mapping $x_{L}:V\rightarrow L$. For a given labeling $x_{L}$ and a node $v\in V$, the view of $v$ consists of the 1-ball $B_{1}(v)$ and the restriction of the port numbering function $\delta$, the input function (if defined) and the labeling $x_{L}$ to $B_{1}(v)$. [^1] The graph induced by the one neighborhood of a node fits within the usual radius 2 verification rules. An example of the view of a node is given in Figure˜1.
Figure 1: Right: A graph with port labels (red) where each node is assigned a label in $\{0,1\}^{*}$ (black). Left: The view of the node marked in gray.
Let $R$ be a set of configurations of the form $(B,x_{L})$ where $B$ is a ball of radius 1 and $x_{L}$ is a labeling defined on the nodes of $B$. We refer to $R$ as a set of local configurations. A configuration is said to be correct with respect to $(L,R)$ if for every $v\in V$ the pair $(B_{1}(v),x_{L}|_{N[v]})$ belongs to $R$, where $x_{L}|_{N[v]}$ denotes the restriction of $x_{L}$ to $N[v]$. In other words, the configuration is correct if the labeling in each ball of radius 1 satisfies the local conditions described by the local configurations in $R$. We refer to the pair $(L,R)$ as a labeling scheme.
A labeling scheme $(L,R)$ is of finite type if for any $1$-ball $B$, there is a finite number of labelings of $B$ in $R$, i.e., the set $|\{x\in C^{V(B)}\mid(B,x)\in R\}|$ is finite. For example, if $L$ is finite then any labeling scheme $(L,R)$ is finite. Observe that a labeling scheme is of finite type if the number of states a node $v$ can take in a correct labeling of its $1$-ball is bounded by a function of its degree $\deg(v)$. Note that in the case of bounded degree graphs, a labeling scheme of finite type corresponds precisely to a classical LCL. In the remainder of this paper we only consider labelings of finite type [^2] A similar concept of Locally Finite Labelings (LFLs) was recently introduced in [schmid2026lfl]..
A labeling scheme is used to characterize the solutions to a given task $\mathcal{T}=(O,P)$. The key property we prove, is that if a configuration is (locally) correct with respect to the labeling scheme, then it is (globally) correct with respect to the task specification (i.e., it is a correct configuration for the task). Formally, there exists $out:L\rightarrow O$, such that there exists a labeling scheme $(L,R)$ for some task $(O,P)$ in a family of graphs (that might initially have inputs) $\mathcal{F}$ if for every $G\in\mathcal{F}$ the following two properties are satisfied [^3] The symbol $\circ$ denotes the composition operator.:
@@
$\forall G\in\mathcal{F},\text{ }\exists(G,x_{L})\text{ s.t. }\forall v\in V,\text{ }(B_{1}(v),x_{L}|_{N[v]})\in R$
@@
$\forall(G,x_{L})\text{ s.t. }G\in\mathcal{F},\text{ if }\forall v\text{ }(B_{1}(v),x_{L}|_{N[v]})\in R\text{ then }(G,\emph{out}\circ x_{L})\in P$
In most of our labeling schemes, the labels correspond to an edge orientation. For these labels, we denote the configuration as $\vec{G}$ and we adopt the following encoding for the labels: the label assigned to each node $v\in V$ is composed of $\deg(v)$ bits where each bit encodes the direction of the corresponding incident edge. We assume that 0 denotes an incoming edge to $v$ and 1 denotes an outgoing edge from $v$. The bit at position $i\in\{0,\ldots,\deg(v)-1\}$ of the label corresponds to the direction of the edge at port $i$ of $v$. A visualization of this for the graph in Figure˜1 is given in Figure˜2.
Figure 2: The $i$-th bit of the label (black numbers) of each node $v$ corresponds to the direction of the edge incident to port $i$ (red numbers) of $v$, for $i\in\{0,\ldots,deg(v)-1\}$ where 0 denotes an incoming and 1 an outgoing edge. On the left we show the direction of the edges and on the right we additionally show the bit of the label of each node corresponding to each edge.